Consider the given shaft where TA and TD are the reaction torques. Screen_Shot_2020-02-27_at_11.37.14_AM.pn

, Part A - Find the ratio between you Part A - Find the ratio between IVO AQ vec* * O O ? Submit Previous Answers Request AnsAB Part C - Determine the ratio- T,L/GJ V AED 1 vecom O2 ? Submit Request Answer ФВС Part D - Determine the ratio TL/GJ IVO APCD Part E - Determine the ratio Lº To L/GJ 187 AX Uvec ec O O ? Submit Request Answer
Answer #1

2 To 10 Torque in AB TAB = I TA A Cap Torque un gc Tec = + TA - TO K S30*4 → Torque en CD :-LAB -Lec = LCD - Teo=TA-To-2 To- 1.5 Tot To = 3 To To = 1.5To To 51.5 Answer To +370 & AB = TAX31 - 1.5 To 3L 5 = GJ AB = 0:45) TOLIGJ Answer (15 To-To) 34